Two solutions are mixed with same kind of solute. Find the Mixed concentration if the concentration of solution 1 is 4.321 %mass with quantity of 1.128 kg and solution 2 has a concentration of 1.161 %mass and has quantity of 1.341 kg?
1
Expert's answer
2021-10-04T05:09:22-0400

m1 = 4.321/100 x 1.128 = 0.049 kg

m2 = 1.161/100 x 1.341 = 0.016 kg


% = (100 x (0.049 + 0.016)) / (1.128 + 1.341) = 2.633%